# Tropical curves of unibranch points and hypertangency

### v1

@@ Abstract.
We study integral plane curves meeting at a single unibranch point and show that such curves must satisfy two equivalent conditions. A numeric condition: the local invariants of the curves at the contact point must be arithmetically related. A geometric condition: the tropical curves that we associate to the contact point must be isomorphic. Moreover, we prove closed formulas for the delta-invariant of a unibranch singularity, and for the dimension of the loci of curves with an assigned unibranch point. Our work is motivated by interest in the Lang exceptional set.
@@
Unibranch singularity, plane curve, hypertangency, euclidean algorithm, tropical curve.
@@ 1.1. Results
In this paper we study unibranch points of curves, i.e. points which have a single preimage in the normalization of the curve, and our main goal is to answer the following question.
Under what conditions can two plane projective curves of degree at least $2$ intersect in exactly one point, $q$, such that $q$ is unibranch for both of them?
This question is also relevant for the study of the Lang exceptional set appearing in some conjectures related to arithmetic and hyperbolic geometry, as we shall explain later in this introduction.
To a unibranch point $q$ of a plane curve $C\subset\mathbb{P}^{2}$ of degree at least $2$ we associate a type made of a pair of integers $(m,n)$, where $m={\operatorname{mult}}_{q}(C)\geq 1$ is the multiplicity of $C$ at $q$, and $n$ is the multiplicity of the intersection of $C$ with its tangent line $L$ at $q$, i.e. $(L\cdot C)_{q}=n\geq 2$; we say that $q$ is a $(m,n)$-point for $C$. The pair $(m,n)$ encodes information on the local geometry of $C$ at $q$. The first main result, answering the above question, numerically describes the connection between the local geometry of two curves meeting at a single unibranch point.
@@
Let $B,C\subset\mathbb{P}^{2}$ be integral curves of degree at least $2$ such that $B\cap C=\{q\}$, where $q$ is an $(m^{\prime},n^{\prime})$-point for $B$ and an $(m,n)$-point for $C$. Then $m^{\prime}/n^{\prime}=m/n$.
The stronger conclusion $(m^{\prime},n^{\prime})=(m,n)$ holds under additional assumptions, see Corollary 4.1.3.
This theorem can be equivalently stated in a more geometric fashion, by associating a tropical curve to a unibranch point of type $(m,n)$. We do that in Section 3 and denote this tropical curve by $\Gamma^{(m,n)}$. Then we prove the following statement, which is an alternative to the previous one
Theorem 1.1.2.
Hypotheses of Theorem 1.1.1. Then there is an isomorphism of tropical curves between $\Gamma^{(m,n)}$ and $\Gamma^{(m^{\prime},n^{\prime})}$.
The core of the proof of these theorems is in Section 4.
Tropical curves are metric graphs, i.e. graphs enriched by the assignment of a length (a positive real number) to each edge; see [MZ08]. In the classical literature non-metric graphs, such as Enriques diagrams, have been widely used to study singularities of plane curves; see [EC85] or [Wal04]. Even though such graphs do not fit our purpose, they inspired our construction of the tropical curve $\Gamma^{(m,n)}$; see Section 3 for more details.
The techniques we develop here enable us to obtain, as a byproduct, closed formulas for the delta-invariant of a unibranch singularity (Proposition 5.1.1) and for the dimension of the space of curves with an assigned unibranch point (Proposition 5.2.1). These formulas occupy the last section of this article.
@@ 1.2. Context
Conjecturally, $\mathcal{E}(B)$ is finite as soon as $\deg B\geq 4$, which corresponds to the pairs $(\mathbb{P}^{2},B)$ being of log general type. More generally, Lang conjectured that the fact that $\mathcal{E}(B)$ is a proper closed subset of $\mathbb{P}^{2}$ should be equivalent to the fact that $(\mathbb{P}^{2},B)$ is of log general type (and this should hold more in general for projective varieties $X$ with a normal crossing divisor $B$). For $\mathbb{P}^{2}$ this conjecture is known only for $B$ reducible with at least three components; the finiteness follows from [CZ13, GNSW23], and we refer to [CT24] for further references and explicit bounds. It remains open when $B$ is irreducible or has only two irreducible components, although [Che04, PR07, CRY23] show that it holds for a very general $B$.
If $C\in\mathcal{E}(B)$ and $C\cap B=\{q,q^{\prime}\}$, then it is clear that $q$ and $q^{\prime}$ must be unibranch points of $C$. Therefore one is naturally led to investigate unibranch points, and curves having high order of contact at such points. In particular when $B$ has (at least) three components, any $C\in\mathcal{E}(B)$ will intersect at least one of the components of $B$ in a single unibranch point. This is precisely the setting of Theorem 1.1.1 and 1.1.2.
As an example of application of this type of results we mention that, in our previous paper [CT24], we used as a crucial ingredient a special case of Theorem 1.1.1 (Theorem 2.2.1 in loc.cit. which assumes $B$ is a smooth at $q$) to show that $\mathcal{E}(B)$ is empty for a general $B$ with three irreducible components of total degree at least 5.
@@ 1.3. Notation
We work over $\mathbb{C}$. Throughout the paper, $S$ will be a smooth projective surface endowed with a birational morphism onto $\mathbb{P}^{2}$, and $C\subset S$ an integral projective curve.
@@ Acknowledgements.
We are grateful to Pietro Corvaja, Kristin DeVleming, Luca Ferrigno and Nicola Ottolini for useful comments. Special thanks to Laura Capuano for showing us Lemma 5.1.2. LC is partially supported by PRIN 2017SSNZAW and PRIN 2022L34E7W, Moduli spaces and birational geometry. AT is partially supported by PRIN 2022HPSNCR: Semiabelian varieties, Galois representations and related Diophantine problems and PRIN 2020KKWT53: Curves, Ricci flat Varieties and their Interactions, and is a member of the INDAM group GNSAGA.
@@ 2.1. Resolving unibranch points
For a unibranch point $q\in C\subset S$ we now introduce an iterated sequence of blow-ups. We set $S_{1}=S$, let $S_{2}\to S_{1}$ be the blow-up at $q$, and $E_{2}\subset S_{2}$ the exceptional divisor. We denote by $C^{2}\subset S_{2}$ the proper transform of $C$ and by $q^{2}\in C^{2}$ the point lying over $q$, i.e. $q^{2}=C^{2}\cap E_{2}$. We use superscripts for strict transforms, and subscripts otherwise, but we denote $C=C^{1}$ and $q=q^{1}$ for convenience. We iterate to get a sequence of arbitrary length
@@
We proceed as in the proof of [CT24, Lemma 2.1.1], of which the present lemma is a generalization. We write, as above, $q^{1}=q\in C^{1}=C\subset S^{1}=S$. We use local coordinates at $q^{1}\in U\cong\mathbb{A}^{2}_{x_{1},y_{1}}$ so that $q^{1}=(0,0)$ and $L=L^{1}$ has equation $y_{1}=0$. As $q^{1}\in C^{1}$ is an $(m,n)$-point w.r.t. $L^{1}$, the defining polynomial, $f_{1}(x_{1},y_{1})$, of $C^{1}$ is
@@
Let $q\in C\subset\mathbb{P}^{2}$ be an $(m,n)$-point, (4) the euclidean sequence of $(m,n)$, and $\nu_{q}:C^{\nu}_{q}\to C$ the desingularization of $C$ at $q$. We have a chain of birational morphisms
@@ 2.2. Normal crossings resolutions
(6) $D_{*}=E_{*}+C_{*}+\sum_{\stackrel{{\scriptstyle j=1,\ldots,r}}{{i=1,\ldots,k_{j}}}}A^{i}_{j}.$
@@ 3.1. Dual graphs of partial resolutions
of the divisor $D_{*}$, defined in Section 2.2. The vertices of $G_{*}$ are the irreducible components of $D_{*}$, hence $G_{*}$ has $2+\sum_{i=1}^{r}k_{i}$ vertices, and there is an edge between two vertices if the two components intersect. For the next lemma, recall that a leaf of a graph is a vertex of degree $1$.
@@ 3.2. The contact tropical curve of an $(m,n)$-point
To prove the converse, write $\Gamma=\Gamma^{(m,n)}$ and $\Gamma^{\prime}=\Gamma^{(m^{\prime},n^{\prime})}$ and assume there is an isomorphism of tropical curves $\phi:\Gamma\to\Gamma^{\prime}$. Hence $\phi$ is a length-preserving isomorphism between the underlying graphs.
@@
Recall that $\Gamma$ and $\Gamma^{\prime}$ have a unique vertex, namely $E_{*}$ and $E^{\prime}_{*}$, of degree $3$, which must, of course, be preserved by an isomorphism. If $r=1$ then $k_{1}\geq 2$, hence $\Gamma$ has a unique leg and looks as follows
@@ 4.1. Hypertangency at unibranch point
The equivalence of the three conditions follows from Remark 3.1.7 and Proposition 3.2.4. We will prove $(m^{\prime},n^{\prime})\sim(m,n)$.
@@
To prove the theorem we used the following Lemma.
@@ 4.2. Intersection multiplicity at equivalent points
We proved in Theorem 4.1.1 that if two curves $B$ and $C$ are hypertangent in one point $q$, then $q\in B$ is equivalent to $q\in C$, i.e. their contact tropical curves are isomorphic. It is natural to pose the opposite problem, namely, what can be said about the intersection of two curves meeting at equivalent points.

### v2

@@ Abstract.
We study plane curves meeting at a single unibranch point and show that they must satisfy two equivalent conditions. A geometric one: the tropical curves associated to the contact point are isomorphic. An arithmetic one: the local invariants at the contact point satisfy a simple relation. Furthermore, we prove closed formulas for the delta-invariant of a unibranch singularity, and for the dimension of the locus of curves with an assigned unibranch point. Our work is motivated by interest in the Lang’s exceptional set.
@@
Unibranch singularity, plane curve, hypertangency, euclidean algorithm, tropical curve, genus.
@@ 1.1. Results
In this paper we study unibranch points of curves, i.e. points which have a single preimage in the normalization of the curve. Our first goal is to study under what conditions two projective plane curves of degree at least $2$ intersect in exactly one point which is unibranch for both of them. This problem is also relevant for the study of the Lang exceptional set appearing in some conjectures related to arithmetic and hyperbolic geometry, as we shall explain later in this introduction.
To a unibranch point $q$ of a plane curve $C\subset\mathbb{P}^{2}$ of degree at least $2$ we associate a type made of a pair of integers $(m,n)$, where $m={\operatorname{mult}}_{q}(C)$ is the multiplicity of $C$ at $q$, and $n$ is the multiplicity of the intersection of $C$ with its tangent line $L$ at $q$, i.e. $(L\cdot C)_{q}=n$; we say that $q$ is a $(m,n)$-point for $C$. Then, in Section 3, we associate a tropical curve, denoted by $\Gamma^{(m,n)}$, to a unibranch point of type $(m,n)$. Our main result answering the above problem can be stated as follows.
@@
Let $B,C\subset\mathbb{P}^{2}$ be integral curves of degree at least $2$ such that $B\cap C=\{q\}$, where $q$ is an $(m^{\prime},n^{\prime})$-point for $B$ and an $(m,n)$-point for $C$. Then there is an isomorphism of tropical curves between $\Gamma^{(m,n)}$ and $\Gamma^{(m^{\prime},n^{\prime})}$; equivalently, $m^{\prime}/n^{\prime}=m/n$.
See Theorem 3.2.4 and Theorem 4.1.1.
Tropical curves are metric graphs, i.e. graphs enriched by the assignment of a length (a positive real number) to each edge; see [MZ08]. In the classical literature non-metric graphs, such as Enriques diagrams, have been widely used to study singularities of plane curves; see [EC85] or [Wal04]. Even though such graphs do not quite fit our purpose, they inspired our construction of the tropical curve $\Gamma^{(m,n)}$; see Subsection 3.1 for more details.
In the second part of Section 4 we investigate the opposite problem, describing the intersection of two curves at points of the same (or, more generally, equivalent) type.
Finally, our techniques enable us to obtain closed formulas for the delta-invariant of a unibranch singularity, in Proposition 5.1.1, and for the dimension of the space of curves with an assigned unibranch point, in Proposition 5.2.1. The proof of these formulas occupy the last section of this article.
@@ 1.2. Context
Conjecturally, $\mathcal{E}(B)$ is finite as soon as $\deg B\geq 4$, which corresponds to the pairs $(\mathbb{P}^{2},B)$ being of log general type. More generally, Lang conjectured that the fact that $\mathcal{E}(B)$ is a proper closed subset of $\mathbb{P}^{2}$ should be equivalent to the fact that $(\mathbb{P}^{2},B)$ is of log general type (and this should hold more in general for projective varieties $X$ with a normal crossing divisor $B$). For $\mathbb{P}^{2}$ this conjecture is known only for $B$ reducible with at least three components; the finiteness follows from [CZ13, GNSW25], and we refer to [CT25] for further references and explicit bounds. It remains open when $B$ is irreducible or has only two irreducible components, although [Che04, PR07, CRY23, ATY24] show that it holds for a very general $B$.
If $C\in\mathcal{E}(B)$ and $C\cap B=\{q,q^{\prime}\}$, then it is clear that $q$ and $q^{\prime}$ must be unibranch points of $C$. Therefore one is naturally led to investigate unibranch points, and curves having high order of contact at such points. In particular when $B$ has (at least) three components, any $C\in\mathcal{E}(B)$ will intersect at least one of the components of $B$ in a single unibranch point. This is precisely the setting of Theorem 1.1.1.
As an example of application of this type of results we mention that, in our previous paper [CT25], we used as a crucial ingredient a special case of Theorem 1.1.1 (Theorem 2.2.1 in loc.cit. which assumes $B$ is a smooth at $q$) to show that $\mathcal{E}(B)$ is empty for a general $B$ with three irreducible components of total degree at least 5.
@@ 1.3. Notation
We work over $\mathbb{C}$. Throughout the paper, $S$ will be a smooth projective surface endowed with a birational morphism to $\mathbb{P}^{2}$, and $C\subset S$ an integral projective curve.
@@ Acknowledgements.
We are grateful to Pietro Corvaja, Kristin DeVleming, Luca Ferrigno, Nicola Ottolini and Wei Chen for useful comments. Special thanks to Laura Capuano for showing us Lemma 5.1.2. LC is partially supported by PRIN 2017SSNZAW and PRIN 2022L34E7W, Moduli spaces and birational geometry. AT is partially supported by PRIN 2022HPSNCR: Semiabelian varieties, Galois representations and related Diophantine problems and PRIN 2020KKWT53: Curves, Ricci flat Varieties and their Interactions, and is a member of the INDAM group GNSAGA.
@@ 2. Unibranch points
This section is devoted to foundational concepts on unibranch points of curves on surfaces.
@@ 2.1. Resolving unibranch points
For a unibranch point $q\in C\subset S$ we introduce an iterated sequence of blow-ups. Set $S_{1}=S$, let $S_{2}\to S_{1}$ be the blow-up at $q$, and $E_{2}\subset S_{2}$ the exceptional divisor. We denote by $C^{2}\subset S_{2}$ the proper transform of $C$ and by $q^{2}\in C^{2}$ the point lying over $q$, i.e. $q^{2}=C^{2}\cap E_{2}$. We use superscripts for strict transforms, and subscripts otherwise, but we denote $C=C^{1}$ and $q=q^{1}$ for convenience. We iterate to get a sequence of arbitrary length
@@
We proceed as in the proof of [CT25, Lemma 2.1.1], of which the present lemma is a generalization. We write, as above, $q^{1}=q\in C^{1}=C\subset S^{1}=S$. We use local coordinates at $q^{1}\in U\cong\mathbb{A}^{2}_{x_{1},y_{1}}$ so that $q^{1}=(0,0)$ and $L=L^{1}$ has equation $y_{1}=0$. As $q^{1}\in C^{1}$ is an $(m,n)$-point w.r.t. $L^{1}$, the defining polynomial, $f_{1}(x_{1},y_{1})$, of $C^{1}$ is
@@
Let $q\in C\subset\mathbb{P}^{2}$ be an $(m,n)$-point, (4) the euclidean sequence of $(m,n)$, and $\nu_{q}:C^{\nu}_{q}\to C$ the desingularization of $C$ at $q$. We have a chain $\beta$ of birational morphisms
@@ 2.2. Normal crossings resolutions
(6) $D_{*}=C_{*}+E_{*}+\sum_{\stackrel{{\scriptstyle j=1,\ldots,r}}{{i=1,\ldots,k_{j}}}}A^{i}_{j}.$
@@ 3. Tropical curves of unibranch points
In this section we introduce the “contact” tropical curve associated to a unibranch point of a curve.
@@ 3.1. Dual graphs of partial resolutions
of the divisor $D_{*}$, defined in Section 2.2. The vertices of $G_{*}$ are the irreducible components of $D_{*}$, hence $G_{*}$ has $2+\sum_{i=1}^{r}k_{i}$ vertices, and there is an edge between two vertices if the two components intersect. For the next lemma, recall that a leaf of a graph is a vertex of degree $1$ (and the degree, or valency, of a vertex is the number of half-edges adjacent to it).
@@ 3.2. The contact tropical curve of an $(m,n)$-point
To prove the converse, write $\Gamma=\Gamma^{(m,n)}$ and $\Gamma^{\prime}=\Gamma^{(m^{\prime},n^{\prime})}$ and assume that there is an isomorphism of tropical curves $\phi:\Gamma\to\Gamma^{\prime}$. Hence $\phi$ is a length-preserving isomorphism between the underlying graphs.
@@
Recall that $\Gamma$ and $\Gamma^{\prime}$ have each a unique vertex, respectively $E_{*}$ and $E^{\prime}_{*}$, of degree $3$. Therefore the isomorphism $\phi$ has to map the vertex $E_{*}$ to the vertex $E^{\prime}_{*}$. If $r=1$ then $k_{1}\geq 2$, hence $\Gamma$ has a unique leg and looks as follows
@@ 4.1. Hypertangency at unibranch point
The proof of the next statement combines our earlier results with results in [GBP15].
@@
The equivalence of the three conditions follows from Remark 3.1.7 and Theorem 3.2.4.
Set $b=\deg B$ and $d=\deg C$. By hypothesis $(B\cdot C)_{q}=(B\cdot C)=bd.$ The assumption implies that $B$ and $C$ have the same tangent line, written $L$, at $q$, so that $(C\cdot L)_{q}=n$ and $(B\cdot L)_{q}=n^{\prime}$.
To prove that $m/n=m^{\prime}/n^{\prime}$ we apply the results in subsection 2.3 of [GBP15], in particular the “Strong Triangle Inequality” in Theorem 2.8 and Corollary 2.9. They imply that the smallest among the following three numbers are equal
(11) $\frac{(B\cdot C)_{q}}{{\operatorname{mult}}_{q}(B){\operatorname{mult}}_{q}(C)},\quad\ \frac{(B\cdot L)_{q}}{{\operatorname{mult}}_{q}(B){\operatorname{mult}}_{q}(L)},\quad\ \frac{(C\cdot L)_{q}}{{\operatorname{mult}}_{q}(C){\operatorname{mult}}_{q}(L)}.$
We have,
$\frac{(B\cdot C)_{q}}{{\operatorname{mult}}_{q}(B){\operatorname{mult}}_{q}(C)}=\frac{bd}{m^{\prime}m}\geq\frac{bn}{m^{\prime}m}>\frac{m^{\prime}n}{m^{\prime}m}=\frac{n}{m}=\frac{(C\cdot L)_{q}}{{\operatorname{mult}}_{q}(C){\operatorname{mult}}_{q}(L)}$
(using $d\geq n$ and $b>m^{\prime}$ for the two inequalities). Similarly
$\frac{(B\cdot C)_{q}}{{\operatorname{mult}}_{q}(B){\operatorname{mult}}_{q}(C)}>\frac{n^{\prime}}{m^{\prime}}=\frac{(L\cdot B)_{q}}{{\operatorname{mult}}_{q}(L){\operatorname{mult}}_{q}(B)}.$
Hence the second and third number in (11) are the smallest, hence they are equal, hence $n/m=n^{\prime}/m^{\prime}$ and we are done.
We now give a different, self-contained proof which applies the material of the previous sections.
Alternative proof.
@@
We used the following elementary Lemma.
@@ 4.2. Intersection multiplicity at equivalent points
We proved in Theorem 4.1.1 that if two curves $B$ and $C$ are hypertangent in one point $q$, then $q\in B$ is equivalent to $q\in C$, i.e. their contact tropical curves are isomorphic. It is natural to pose the opposite problem, namely, what can be said about the intersection of two curves at equivalent points.